. 0.15 m3 of an ideal gas at a pressure of 15 bar and 550 K is expanded
isothermally to 4 times the initial volume. It is then cooled to 290 K at constant volume and then
compressed back polytropically to its initial state.
Calculate the net work done and heat transferred during the cycle.
1
Expert's answer
2021-09-10T12:17:23-0400

Given :

"V_1 = 0.15 \;m^3 \\

 p_1 = 15 \; bar \\

 T_1 = T_2 = 550\; K \\

\frac{ V_2}{V_1} = 4 \\

 T_3 = 290 \; K"



Considering the isothermal process 1–2, we have

"p_1 V_1 = p_2 V_2 \\

 p_2 = \frac{p_1V_1}{V_2} \\

p_2 = \frac{15 \times 0.15}{4 \times 0.15} = 3.75 \;bar"

Work done,

"W_{1-2} = p_1V_1 log_e(\frac{V_2}{V_1}) \\

= (15 \times 10^5 ) \times 0.15 \times log_e (4) \\

= 311916 \; J = 311.9 \; kJ"

Considering constant volume process 2-3, we get

"V_2=V_3=V_4 \times 0.15 =0.6 \; m^3 \\

\frac{p_2}{T_2} = \frac{p_3}{T_3} \\

p_3=p_2 \times \frac{T_3}{T_2} = 3.75 \times \frac{290}{550} = 1.98 \;bar \\

W_{2-3} =0"

...since volume remains constant

Consider polytropic process 3–1 :

"p_3V_3^n = p_1V_1^n \\

\frac{p_1}{p_3} = (\frac{V_3}{V_1})^n"

Taking log on both sides, we get

"log_e(\frac{p_1}{p_3}) = n log_e( \frac{V_3}{V_1}) \\

n = \frac{log_e(\frac{p_1}{p_3})}{log_e( \frac{V_3}{V_1})} \\

= \frac{log_e(\frac{15}{1.98})}{log_e(4)} \\

= 1.46 \\

W_{3-1} = \frac{p_3V_3 -p_1V_1}{n-1} \\

= \frac{1.98 \times 10^5 \times 0.6 – 15 \times 10^5 \times 0.15}{1.46-1} \\

= -230869 \;J \\

= -230.87 \;kJ"

Net work done "= W_{1\u20132} + W_{2\u20133} + W_{3\u20131}"

"= 311.9 + 0 +(-230.87) = 81.03 \;kJ"

For a cuclic process,

"\oint δQ= \oint δW"

Heat transferred during the cycle = 81.03 kJ